If the consumption function is







C = 50+2√Y







calculate MPC and MPS when Y = 36 and give an interpretation of these results.

1
Expert's answer
2021-11-09T17:52:58-0500

According to the guidelines only one question will be answered please upload other question separately.

(a)

MPC is determined as the slope of the consumption function.

"MPC=\frac{dC}{dy}\\=\frac{d}{dy}(50+2\sqrt y)\\=y^{\frac{−1}{2}}=36^{\frac{-1}{2}}=0.1667"


To determine MPS at y=36 substitute 36 for y in MPC.


"MPS=1−(36)^{−\frac{1}{2}}\\=1−\frac{1}{6}\\=\frac{5}{6}=0.833"